# Algebraic and geometric properties of homeomorphism groups of ordinals

### v1

@@ 2.4 Topology of homeomorphism groups
Though our focus is on homeomorphism groups as abstract groups, a key subgroup in what follows is defined by taking a closure in an appropriate topology on the homeomorphism group of an ordinal.
@@
For any ordinal $\alpha$, the group $\operatorname{Homeo}(\alpha+1)$ equipped with the permutation topology is a topological group. ∎
@@ 3 Cantor–Bendixson degree one
Throughout this section, we study $\operatorname{Homeo}(\omega^{\alpha+1})$ instead of $\operatorname{Homeo}(\omega^{\alpha+1}+1)$, but recall they are isomorphic by 2.21.
Before beginning, let us recall two standard definitions. Let $X$ be a topological space. A family of subset $\{Y_{n}\}_{n\in\mathbb{N}}$ of $X$ is locally finite if, given any compact set $K$, the cardinality of the set $\{n\in\mathbb{N}:K\cap Y_{n}\neq\varnothing\}$ is finite. The support of a homeomorphism $f\colon\thinspace X\to X$ is the closure of the set $\{x\in X:f(x)\neq x\}$. Given a sequence $\{f_{n}\}_{n\in\mathbb{N}}$ of homeomorphisms $X\to X$ whose supports form a locally finite family of pairwise-disjoint sets, the infinite product of the $f_{n}$, denoted $\prod_{n\in\mathbb{N}}f_{n}$, is the homeomorphism that agrees with $f_{n}$ on the support of $f_{n}$ for each $n\in\mathbb{N}$ and restricts to the identity elsewhere. In the compact-open topology on $\operatorname{Homeo}(X)$, the infinite product can be realized as the limit of the finite products, that is, $\prod_{n\in\mathbb{N}}f_{n}=\lim_{n\to\infty}\prod_{k=1}^{n}f_{k}$.
@@ 3.1 Topological moieties
Let $A$ be a moiety of $\omega^{\alpha+1}$. Then $A$, equipped with the order inherited from $\omega^{\alpha+1}$, is a well-ordered set, and hence order isomorphic to an ordinal. Let $\{a_{n}\}_{n\in\mathbb{N}}$ be an enumeration of the maximal rank elements of $A$ such that $a_{n}<a_{m}$ whenever $n<m$. Let $U_{1}=[0,a_{1}]\cap A$, and let $U_{n}=[a_{n-1}+1,a_{n}]\cap A$ for $n>1$. Then $\{U_{n}\}_{n\in\mathbb{N}}$ is a collection of pairwise-disjoint clopen subsets satisfying $A=\bigcup_{n\in\mathbb{N}}U_{n}$, implying that $A$ is homeomorphic to the topological disjoint union of the $A_{n}$. Now, by the classification of successor ordinals (2.15), each of the $U_{n}$ is homeomorphic to $\omega^{\alpha}+1$; therefore, $A$ is homeomorphic to ×$\mathbb{N}\times(\omega^{\alpha}+1)$, which is in turn homeomorphic to $\omega^{\alpha+1}$ by 2.22. ∎
@@
Let $\alpha$ be an ordinal. If $A$ and $B$ are topological moieties, then there exists $\sigma\in\operatorname{Homeo}(\omega^{\alpha+1})$ such that $\sigma(A)=B$. Moreover, if $A\cap B=\varnothing$, then $\sigma$ can be chosen to be an involution supported in $A\cup B$.
@@
By 3.2, there exists a homeomorphism $f\colon\thinspace A\to B$. Let $C$ and $D$ the complements of $A$ and $B$, respectively. And as the complement of a topological moiety is a topological moiety, there exists a homeomorphism $g\colon\thinspace C\to D$. Define $\sigma\in\operatorname{Homeo}(\omega^{\alpha+1})$ such that $\sigma|_{A}=f$ and $\sigma|_{C}=g$. Then $\sigma(A)=B$.
@@
By 2.22, $\omega^{\alpha+1}$ is homeomorphic to ×$\mathbb{N}\times(\omega^{\alpha}+1)$, which in turn is homeomorphic to ×$\mathbb{Z}\times\mathbb{N}^{2}\times(\omega^{\alpha}+1)$. Consider the subset $A^{\prime}$ of the latter space given by $A^{\prime}=\{0\}\times\{1\}\times\mathbb{N}\times(\omega^{\alpha}+1)$. Define $\tau^{\prime}$ by $\tau^{\prime}(\ell,1,n,x)=(\ell+1,1,n,x)$ and $\tau^{\prime}(\ell,m,n,x)=(\ell,m,n,x)$ whenever $m>1$. Fix a homeomorphism $\psi\colon\thinspace\mathbb{Z}\times\mathbb{N}^{2}\times(\omega^{\alpha}+1)\to\omega^{\alpha+1}$. Then $\psi(A^{\prime})$ is a topological moiety, and by 3.3, we can assume that $\psi(A^{\prime})=A$. It follows that $\varphi=\psi\circ\tau^{\prime}\circ\psi^{-1}$ is a convergent $A$-translation. Moreover, as $\tau^{\prime}$ is supported in ×$\mathbb{Z}\times\{1\}\times\mathbb{N}\times(\omega^{\alpha}+1)$, we have that $\tau$ is supported in a topological moiety. ∎
@@ 3.2 Galvin’s lemma
The key behind all the proofs in the remainder of the section is a uniform fragmentation result, i.e., given two topological moieties whose union is a topological moiety, we provide a way of writing any homeomorphism as a composition of three homeomorphisms each of which is supported in one of the two given moieties. We call this uniform fragmentation result Galvin’s lemma, as it is an extension of Galvin’s [14, Lemma 2.1] from the setting of permutation groups to $\operatorname{Homeo}(\omega^{\alpha+1})$. Given a topological moiety $A$ of $\omega^{\alpha+1}$, we let
@@
Fix $h\in\operatorname{Homeo}(\omega^{\alpha+1})$. Let $C$ denote the complement of $A\cup B$. As
@@ 3.3 Normal generation and uniform perfectness
Let us recall several definitions. A group $G$ is perfect if it is equal to its commutator subgroup, i.e., the subgroup generated by the set $\{[x,y]:x,y\in G\}$, where $[x,y]=xyx^{-1}y^{-1}$; it is uniformly perfect if there exists some $k\in\mathbb{N}$ such that every element of $G$ can be expressed as a product of $k$ commutators, and the minimum such $k$ is called the commutator width of $G$.
@@
Let $A$ be a topological moiety that contains the support of $h$. By 3.5, there exists a convergent $A$-translation $\tau\in\operatorname{Homeo}(\omega^{\alpha+1})$. We now apply a standard commutator trick. Setting $\sigma=\prod_{n=0}^{\infty}(\tau^{n}\circ h\circ\tau^{-n})$, it is readily checked that $h=[\sigma,\tau]$. ∎
@@
Informally, one sees this by noting that $\sigma$ is “performing $f$ on all the forward iterates of $B$ in A”, and hence so is $\eta$, but $\eta$ is also “performing $f^{-1}$ on all the forward iterates of $h(B)$ in $h(A)$”. Now, $\psi\circ\eta\circ\psi^{-1}$ restricts to the inverse of $\eta$ on the complement of $B$ and is the identity on $B$ (this is on account of $\psi|_{h(A)}=\tau\circ h^{-1}|_{h(A)}$). Therefore, the composition $\eta\circ(\psi\circ\eta\circ\psi^{-1})$ is supported on $B$, where it agrees with $\sigma$ and hence with $f$.
@@ 3.4 Strong distortion
Recall that a group $G$ is strongly distorted if there exists $m\in\mathbb{N}$ and a sequence $\{w_{n}\}_{n\in\mathbb{N}}\subset\mathbb{N}$ such that for any sequence $\{g_{n}\}_{n\in\mathbb{N}}\subset G$ there exists $S\subset G$ of cardinality at most $m$ such that $g_{n}\in S^{w_{n}}$. The notion of strong distortion was introduced by Calegari–Freedman [9], and they provide a technique for establishing strong distortion that we employ here. This technique was described in a general form by Le Roux–Mann [21, Construction 2.3] that we rephrase here in the context of ordinals. Below, and throughout the rest of the article, given $g,h\in G$, we let $g^{h}=hgh^{-1}$.
Proposition 3.13.
@@
Let $\{h_{n}\}_{n\in\mathbb{N}}$ be a sequence in $\operatorname{Homeo}(\omega^{\alpha+1})$. We must find $m\in\mathbb{N}$ and a sequence $\{w_{n}\}_{n\in\mathbb{N}}$ in $\mathbb{N}$, independent of $\{h_{n}\}$, such that $h_{n}$ is a word of length at most $w_{n}$ in a subgroup generated by a subset $S$, depending on $\{h_{n}\}$, of $\operatorname{Homeo}(\omega^{\alpha+1})$ of cardinality at most $m$.
Fix a topological moiety $A$. Let us first assume that each of the $h_{n}$ is supported in $A$. By 3.5, there exists a convergent $A$-translation $\sigma$ supported in a topological moiety $B$. Using 3.5 again, there exists a convergent $B$-translation $\tau$. Applying 3.13 yields a homeomorphism $\varphi$ such that $h_{n}$ can be expressed as a word of length $44n+$ 4 in the set $\{\varphi,\sigma,\tau\}$.
We now assume $\{h_{n}\}$ is a general sequence, but we keep $A$, $B$, $\sigma$, $\tau$, and $\varphi$ as above. Fix a topological moiety $C$ in $\omega^{\alpha+1}$ such that $A\cup C$ is a topological moiety, and using 3.3, choose $\theta\in\operatorname{Homeo}(\omega^{\alpha+1})$ such that $\theta(C)=A$. By 3.6, we can write
@@
where either $h_{n,i}$ or $h_{n,i}^{\theta}$ is supported in $A$. Therefore, by the above argument, we can write $h_{n}$ as a word of length at most $1812n+$ 18 in the subgroup generated by the set $S=\{\sigma,\tau,\varphi,\theta\}$. Setting $m=4$ and $w_{n}=12n+18$ establishes the result. ∎
@@ 4 Cantor–Bendixson degree greater than one
4.1 Right-split short exact sequences
The goal of this subsection is to establish two right-split exact sequences, one of the form
@@ 4.1 Right-split short exact sequences
Let $G_{k}$ denote the subgroup of $\operatorname{PH}$ consisting of elements supported in $\Omega_{k}$. Note that $f_{k}\in G_{k}$. We can identify $G_{k}$ with $\operatorname{Homeo}(\Omega_{k})$, and as $\Omega_{k}$ is homeomorphic to $\omega^{\alpha+1}+1$ by 2.18, we can apply 1.1 to see that $G_{k}\cap\overline{F}=G_{k}$, allowing us to conclude that $f_{k}\in\overline{F}$, as desired. ∎
@@
[…]
    is a short exact sequence and it is right split.
@@
    If $\Phi\colon\thinspace\operatorname{H}_{\alpha,d}\to\operatorname{Sym}(M_{\alpha,d})$ is the canonical action, then the short exact sequence
    $1\longrightarrow\operatorname{PH}_{\alpha,d}\lhook\joinrel\longrightarrow\operatorname{H}_{\alpha,d}\overset{\Phi}{\longrightarrow}\operatorname{Sym}(M_{\alpha,d})\longrightarrow 1$
    is right split.
@@
For (1), we have already established the section to $\chi$ in 4.1; in 4.2, we showed that $\overline{\operatorname{F}}$ is in $\ker\chi$. It is left to show that $\ker\chi$ is in $\overline{\operatorname{F}}$. Let $f\in\ker\chi$. Then 4.3 implies $f=f_{0}\circ f_{1}\circ\cdots\circ f_{k}$ with $f_{0}\in\operatorname{F}$ and $f_{k}$ supported in $\Omega_{k}$. Let $G_{k}$ denote the subgroup of $\operatorname{PH}$ consisting of elements supported in $\Omega_{k}$. We can identify $G_{k}$ with $\operatorname{Homeo}(\Omega_{k})$, and as $\Omega_{k}$ is homeomorphic to $\omega^{\alpha+1}+1$ by 2.18, we can apply 1.1 to see that $G_{k}\cap\overline{\operatorname{F}}=G_{k}$, allowing us to conclude that $f_{k}\in\overline{\operatorname{F}}$, as desired. This establishes (1).
We move to the second sequence, which is exact by definition; we must show that it is right split. Let $J_{1}=\{x\in X_{\alpha,d}:x\leq\mu_{1}\}$ and, for $k\in\{2,\ldots,d\}$, let $J_{k}=\{x\in X_{\alpha,d}:\mu_{k-1}<x\leq\mu_{k}\}$. Given $j,k\in\{1,\ldots,d\}$, there is a unique order-preserving bijection $g_{j,k}\colon\thinspace J_{j}\to J_{k}$. Using our enumeration of $M$, identify $\operatorname{Sym}(M)$ with $\operatorname{Sym}\left(\{1,\ldots,d\}\right)$. Taking $\sigma\in\operatorname{Sym}(M)$, define $\Psi(\sigma)\in\operatorname{H}$ to be the homeomorphism given by
@@
whenever $x\in J_{k}$. Then $\Psi$ is a section of $\Phi$, finishing the proof. ∎
Before continuing, let us come back to noting that the definition of $\chi$ did not depend on the choice of the $\Omega_{k}$. Suppose a different choice is made resulting in a homomorphism $\chi^{\prime}$. A careful reading of the proof of Lemma 4.1 shows that $\chi$ and $\chi^{\prime}$ share a section. Moreover, Theorem 4.4 applies to $\chi^{\prime}$, so that $\chi$ and $\chi^{\prime}$ have the same kernel. Together, these facts imply that $\chi=\chi^{\prime}$.

### v2

@@ 2.4 Topology of homeomorphism groups
Though our main focus is on homeomorphism groups as abstract groups, a key subgroup in what follows is defined by taking a closure in an appropriate topology on the homeomorphism group of an ordinal; moreover, 4.4 is strengthened by making a statement in the category of topological groups, rather than that of abstract groups.
Let us recall that a topological group $G$ is an abstract group equipped with a Hausdorff topology such that the multiplication map $G\times G\to G$ given by $(g,h)\mapsto gh$ and the inversion map $G\to G$ given by $g\mapsto g^{-1}$ are continuous.
@@
For any ordinal $\alpha$, the permutation topology and the compact-open topology on $\operatorname{Homeo}(\alpha+1)$ agree and $\operatorname{Homeo}(\alpha+1)$ equipped with the compact-open topology, equivalently the permutation topology, is a topological group. ∎
@@ 3 Cantor–Bendixson degree one
Throughout this section, we study $\operatorname{Homeo}(\omega^{\alpha+1})$ instead of $\operatorname{Homeo}(\omega^{\alpha+1}+1)$, but recall they are isomorphic by 2.24.
Before beginning, let us recall two standard definitions. Let $X$ be a topological space. A family of subsets $\{Y_{n}\}_{n\in\mathbb{N}}$ of $X$ is locally finite if, given any compact set $K$, the cardinality of the set $\{n\in\mathbb{N}:K\cap Y_{n}\neq\varnothing\}$ is finite. The support of a homeomorphism $f\colon\thinspace X\to X$ is the closure of the set $\{x\in X:f(x)\neq x\}$. Given a sequence $\{f_{n}\}_{n\in\mathbb{N}}$ of homeomorphisms $X\to X$ whose supports form a locally finite family of pairwise-disjoint sets, the infinite product of the $f_{n}$, denoted $\prod_{n\in\mathbb{N}}f_{n}$, is the homeomorphism that agrees with $f_{n}$ on the support of $f_{n}$ for each $n\in\mathbb{N}$ and restricts to the identity elsewhere. In the compact-open topology on $\operatorname{Homeo}(X)$, the infinite product can be realized as the limit of the finite products, that is, $\prod_{n\in\mathbb{N}}f_{n}=\lim_{n\to\infty}\prod_{k=1}^{n}f_{k}$.
@@ 3.1 Topological moieties
Let $A$ be a moiety of $\omega^{\alpha+1}$. Then $A$, equipped with the order inherited from $\omega^{\alpha+1}$, is a well-ordered set, and hence order isomorphic to an ordinal. Let $\{a_{n}\}_{n\in\mathbb{N}}$ be an enumeration of the maximal rank elements of $A$ such that $a_{n}<a_{m}$ whenever $n<m$. Let $U_{1}=[0,a_{1}]\cap A$, and let $U_{n}=[a_{n-1}+1,a_{n}]\cap A$ for $n>1$. Then $\{U_{n}\}_{n\in\mathbb{N}}$ is a collection of pairwise-disjoint clopen subsets satisfying $A=\bigcup_{n\in\mathbb{N}}U_{n}$, implying that $A$ is homeomorphic to the topological disjoint union of the $A_{n}$. Now, by the classification of successor ordinals (2.18), each of the $U_{n}$ is homeomorphic to $\omega^{\alpha}+1$; therefore, $A$ is homeomorphic to $\mathbb{N}\times(\omega^{\alpha}+1)$, which is in turn homeomorphic to $\omega^{\alpha+1}$ by 2.25. ∎
@@
Let $\alpha$ be an ordinal. If $A$ and $B$ are topological moieties, then there exists $\sigma$ in $\operatorname{Homeo}(\omega^{\alpha+1})$ such that $\sigma(A)=B$. Moreover, if $A\cap B=\varnothing$, then $\sigma$ can be chosen to be an involution supported in $A\cup B$.
@@
By 3.2, there exists a homeomorphism $f\colon\thinspace A\to B$. Let $C$ and $D$ the complements of $A$ and $B$, respectively. As the complement of a topological moiety is a topological moiety, there exists a homeomorphism $g\colon\thinspace C\to D$. Define $\sigma\in\operatorname{Homeo}(\omega^{\alpha+1})$ such that $\sigma|_{A}=f$ and $\sigma|_{C}=g$. Then $\sigma(A)=B$.
@@
By 2.25, $\omega^{\alpha+1}$ is homeomorphic to $\mathbb{N}\times(\omega^{\alpha}+1)$, which in turn is homeomorphic to $\mathbb{Z}\times\mathbb{N}^{2}\times(\omega^{\alpha}+1)$. Consider the subset $A^{\prime}$ of the latter space given by $A^{\prime}=\{0\}\times\{1\}\times\mathbb{N}\times(\omega^{\alpha}+1)$. Define $\tau^{\prime}$ by $\tau^{\prime}(\ell,1,n,x)=(\ell+1,1,n,x)$ and $\tau^{\prime}(\ell,m,n,x)=(\ell,m,n,x)$ whenever $m>1$. Fix a homeomorphism $\psi\colon\thinspace\mathbb{Z}\times\mathbb{N}^{2}\times(\omega^{\alpha}+1)\to\omega^{\alpha+1}$. Then $\psi(A^{\prime})$ is a topological moiety, and by 3.3, we can assume that $\psi(A^{\prime})=A$. It follows that $\varphi=\psi\circ\tau^{\prime}\circ\psi^{-1}$ is a convergent $A$-translation. Moreover, as $\tau^{\prime}$ is supported in $\mathbb{Z}\times\{1\}\times\mathbb{N}\times(\omega^{\alpha}+1)$, we have that $\tau$ is supported in a topological moiety. ∎
@@ 3.2 Galvin’s lemma
The key behind all the proofs in the remainder of the section is a uniform fragmentation result, i.e., given two topological moieties whose union is a topological moiety, we provide a way of writing any homeomorphism as a composition of three homeomorphisms each of which is supported in one of the two given moieties. We call this uniform fragmentation result Galvin’s lemma, as it is an extension of Galvin’s [12, Lemma 2.1] from the setting of permutation groups to $\operatorname{Homeo}(\omega^{\alpha+1})$. Given a topological moiety $A$ of $\omega^{\alpha+1}$, we let
@@
Fix $h\in\operatorname{Homeo}(\omega^{\alpha+1})$. Let $C$ denote the complement of $A\sqcup B$. As
@@ 3.3 Normal generation and uniform perfectness
Let us recall several definitions. A group $G$ is perfect if it is equal to its commutator subgroup, i.e., the subgroup generated by the set $\{[x,y]:x,y\in G\}$, where $x,y]=xyx^{-1}y^{-1}$; it is uniformly perfect if there exists some $k\in\mathbb{N}$ such that every element of $G$ can be expressed as a product of $k$ commutators, and the minimum such $k$ is called the commutator width of $G$.
@@
Let $A$ be a topological moiety that contains the support of $h$. By 3.5, there exists a convergent $A$-translation $\tau\in\operatorname{Homeo}(\omega^{\alpha+1})$. We now apply a standard commutator trick. Observe that the supports of the homeomorphisms $\tau^{n}\circ h\circ\tau^{-n}$ form a locally finite family of pairwise-disjoint sets, allowing us to define the infinite product $\sigma=\prod_{n=0}^{\infty}(\tau^{n}\circ h\circ\tau^{-n})$. It can now readily be checked that $h=[\sigma,\tau]$. ∎
@@
Informally, one sees this by noting that $\sigma$ is “performing $f$ on all the forward iterates of $B$ in A”, and hence so is $\eta$, but $\eta$ is also “performing $f^{-1}$ on all the forward iterates of $h(B)$ in h(A)h(A)”. Now, $\psi\circ\eta\circ\psi^{-1}$ restricts to the inverse of $\eta$ on the complement of $B$ and is the identity on $B$ (this is on account of $\psi|_{h(A)}=\tau\circ h^{-1}|_{h(A)}$). Therefore, the composition $\eta\circ(\psi\circ\eta\circ\psi^{-1})$ is supported on $B$, where it agrees with $\sigma$ and hence with $f$.
@@ 3.4 Strong distortion
A group $G$ is strongly distorted if there exists $m\in\mathbb{N}$ and a sequence $\{w_{n}\}_{n\in\mathbb{N}}\subset\mathbb{N}$ such that for any sequence $\{g_{n}\}_{n\in\mathbb{N}}\subset G$ there exists $S\subset G$ of cardinality at most $m$ such that $g_{n}\in S^{w_{n}}$. Above, for $n\in\mathbb{N}$, $S^{n}=\{s_{1}s_{2}\cdots s_{n}:s_{i}\in S\}$ consists of the elements in $G$ that can be written as a word of length $n$ in $S$. The notion of strong distortion was introduced by Calegari–Freedman [8], and they provide a technique for establishing strong distortion that we employ here. This technique was described in a general form by Le Roux–Mann [19, Construction 2.3] that we rephrase here in the context of ordinals. Below, and throughout the rest of the article, given $g,h\in G$, we let $g^{h}=h^{-1}gh$.
Proposition 3.13 ([19, Construction 2.3] for ordinals).
@@
Let $w_{n}=12n+18$ for $n\in\mathbb{N}$. We will show that given any sequence $\{h_{n}\}_{n\in\mathbb{N}}\subset\operatorname{Homeo}(\omega^{\alpha+1})$, there exists $S\subset\operatorname{Homeo}(\omega^{\alpha+1})$ of cardinality four satisfying $h_{n}\in S^{w_{n}}$, thereby establishing that $\operatorname{Homeo}(\omega^{\alpha+1})$ is strongly distorted.
Fix a topological moiety $A$. Let us first assume that each of the $h_{n}$ is supported in $A$. By 3.5, there exists a convergent $A$-translation $\sigma$ supported in a topological moiety $B$. Using 3.5 again, there exists a convergent $B$-translation $\tau$. Applying 3.13 yields a homeomorphism $\varphi$ such that $h_{n}$ can be expressed as a word of length $4n+4$ in the set $\{\varphi,\sigma,\tau\}$.
We now assume $\{h_{n}\}$ is a general sequence, but we keep $A$, $B$, $\sigma$, $\tau$, and $\varphi$ as above. Fix a topological moiety $C$ in $\omega^{\alpha+1}$ such that $A\cup C$ is a topological moiety, and using 3.3, choose $\theta\in\operatorname{Homeo}(\omega^{\alpha+1})$ such that $\theta(A)=C$. By 3.6, we can write
@@
where either $h_{n,i}$ or $h_{n,i}^{\theta}$ is supported in $A$. Therefore, by the above argument, we can write $h_{n}$ as a word of length $12n+18$ in the subgroup generated by the set $S=\{\sigma,\tau,\varphi,\theta\}$, as desired. ∎
@@ 4 Cantor–Bendixson degree greater than one
4.1 Split short exact sequences
The goal of this subsection is to establish two split exact sequences, one of the form
@@ 4.1 Split short exact sequences
In particular, $\ker\chi_{\alpha,d}$ is contained in $\overline{\operatorname{F}}_{\alpha,d}$.
@@
Let $G_{k}$ denote the subgroup of $\operatorname{PH}$ consisting of elements supported in $\Omega_{k}$. Note that $f_{k}\in G_{k}$. We can identify $G_{k}$ with $\operatorname{Homeo}(\Omega_{k})$, and as $\Omega_{k}$ is homeomorphic to $\omega^{\alpha+1}+1$ by 2.21, we can apply 1.1 to see that $G_{k}\cap\overline{F}=G_{k}$, allowing us to conclude that $f_{k}\in\overline{F}$, as desired. ∎
@@
    $1\longrightarrow\overline{\operatorname{F}}_{\alpha,d}\lhook\joinrel\longrightarrow\operatorname{PH}_{\alpha,d}\overset{\chi}{\longrightarrow}\mathbb{Z}^{d-1}\longrightarrow 1$
    is a split short exact sequence.
@@
    The short exact sequence
    $1\longrightarrow\operatorname{PH}_{\alpha,d}\lhook\joinrel\longrightarrow\operatorname{H}_{\alpha,d}\longrightarrow\operatorname{Sym}(d)\longrightarrow 1$
    is split.
Moreover, all the homomorphisms—including the sections—in the above exact sequences are continuous when $\operatorname{H}_{\alpha,d}$, $\overline{\operatorname{F}}_{\alpha,d}$, and $\operatorname{PH}_{\alpha,d}$ are equipped with the compact-open topology and $\mathbb{Z}^{d-1}$ and $\operatorname{Sym}(d)$ are equipped with the discrete topology.
@@
For (1), we have already established the section to $\chi$ in 4.1; in 4.2, we showed that $\overline{\operatorname{F}}$ is contained in $\ker\chi$; and in 4.3, we showed that $\ker\chi$ is contained in $\overline{F}$. We have therefore established (1).
We move to the second sequence, which is exact by definition; we must show that it is split. Let $J_{1}=\{x\in X_{\alpha,d}:x\leq\mu_{1}\}$ and, for $k\in\{2,\ldots,d\}$, let $J_{k}=\{x\in X_{\alpha,d}:\mu_{k-1}<x\leq\mu_{k}\}$. Given $j,k\in\{1,\ldots,d\}$, there is a unique order-preserving bijection $g_{j,k}\colon\thinspace J_{j}\to J_{k}$. Using our enumeration of $M$, identify $\operatorname{Sym}(M)$ with $\operatorname{Sym}(d)$, and let $\Phi\colon\thinspace\operatorname{H}\to\operatorname{Sym}(d)$ be the homomorphism corresponding to the action of $\operatorname{H}$ on $M$, so $\ker\Phi=\operatorname{PH}$. Taking $\sigma\in\operatorname{Sym}(M)$, define $\Psi(\sigma)\in\operatorname{H}$ to be the homeomorphism given by
@@
whenever $x\in J_{k}$. Then $\Psi$ is a section of $\Phi$, establishing (2).
Now to finish, we must establish the continuity of the homomorphisms in (1) and (2). The inclusions are continuous by definition; therefore, it is only left to verify the continuity of $\chi$ and $\Phi$. Recall that (i) a homomorphism from a topological group to a discrete group is continuous if and only if the kernel is open, and (ii) a subgroup of a topological group is open if and only if it contains a neighborhood of the identity. By (1), the kernel of $\chi$ is $\overline{\operatorname{F}}$, which is readily verified to be open. Indeed, $\overline{\operatorname{F}}$ contains the open neighborhood of the identity containing the elements that stabilize each of the $\Omega_{k}$ setwise. And, as $\operatorname{PH}$ contains $\overline{\operatorname{F}}$, which is open, $\operatorname{PH}$ is open itself, implying $\Phi$ is continuous. ∎
Before continuing, let us come back to noting that the definition of $\chi$ did not depend on the choice of the $\Omega_{k}$. Indeed, let $\Omega_{1}^{\prime},\ldots,\Omega_{k}^{\prime}$ be another such partition and let $\chi^{\prime}$ be the resulting homomorphism. Then there exists $h\in\operatorname{PH}$ such that $h(\Omega_{k})=\Omega_{k}^{\prime}$ for all $k$. It follows that $\chi^{\prime}=\chi\circ i_{h}$, where $i_{h}$ denotes conjugation by $h$. As the codomain is abelian, we deduce that $\chi^{\prime}=\chi$.